# Are language models rational? The case of coherence norms and belief revision

### v1

@@ 1 Introduction
To get clearer on our main topic, it is important to first note that the term ‘rational’ is used in different ways, and depending on how one understands it, our main question either becomes a trivial one or a more substantial one. On the one hand, ‘rational’ is often used simply as a term of approval, where an AI is rational just in case it does what it is supposed to do. See Russell, (1997). On such an understanding, intelligent systems are rational, as long as they work properly. But this conception of being rational does not consider how the system does what it is supposed to do, and what internal states of the system contribute to its success. A different sense of ‘rational’ focuses on these internal states instead, not just on the overall behavior of the system. This is how we often evaluate human beings as rational. Here, we do not simply evaluate their behavior, but we focus on their internal states and evaluate them. We consider their beliefs or reasoning or actions as rational, and we thereby evaluate them as having arisen in the proper way from the proper internal, representational states. We can call the former, thinner notion of being rational the deflationary notion, and the latter the representational notion, since it concerns internal representational states.
@@
In the literature on language models, there is much discussion about whether language models have the ability to reason, which is a cross-temporal rational ability. See Huang and Chang, (2023). Reasoning involves a sequence of steps that are properly rationally supported by earlier stages in the process. But the question whether language models engage in reasoning is hard to properly address without first settling whether language models are rational at all and whether norms of rationality even apply to them. The simplest case for making progress on that basic question is to focus first on synchronic coherence norms and on belief states governed by such norms.
Thus to ask whether language models are subject to norms of (representational) rationality like coherence norms involves asking firstly whether they have the proper internal representational states like belief states that are in principle subject to such norms, and secondly whether the models with those states indeed are subject to such norms. To get clear on this issue it is important to distinguish two ways in which a language model can fail to be rational. The first concerns whether a language model is under the norms of rationality in the first place. Here ‘rational’ contrasts with being arational: not being subject to the norms of rationality. The second concerns whether the model lives up to the norms of rationality that apply to it. Here ‘rational’ contrasts with being irrational: not living up the norms of rationality that apply. A toaster is not rational in the first sense, while a person with contradictory beliefs is not rational in the second sense. An arational toaster does not fail to live up the standards of rationality, since those standards do not apply to it at all, while an irrational person does fail this way. We need to ask both of these questions of language models: do the norms of rationality apply to them at all, and if so, do the models live up to those norms. In this paper we will focus mostly on the first, more normative or philosophical question.
@@ 2 Belief in language models
The easiest way to argue that language models have beliefs is via the connection of belief to knowledge. Language models seem to know a lot about the world. Although they might make things up and get things wrong, they nonetheless often reliably produce correct answers to many questions about the world. For example, many large language models can correctly answer the question of what the capital of France is and thus are naturally taken to know that the capital of France is Paris. But knowledge is just a special kind of belief: a belief that is true and especially well supported. The details of the connection between knowledge and belief have long been controversial in the philosophical literature. In particular, it is controversial what else is required besides a state being a belief state that has a true content for this state to be knowledge. Nonetheless, the general connection between knowledge and belief relied upon here is almost universally accepted: to know is to believe. Thus since language models know about the world, they have beliefs about the world. [^1] For a discussion of the relationship between belief and knowledge, see Ichikawa and Steup, (2018).
@@
To have beliefs about the world requires an internal state that is a state of believing, and simply carrying information alone is not enough for this. The question is whether language models have internal states which are states of belief. There are two main worries about this, one is very familiar and one less so. The first concerns whether the internal states of language models are representational states at all. Do any internal states of the model represent the world which created the text on which they are trained? Or does the model merely reliably predict text without any regard to the semantic properties of the text they produce? This issue is widely discussed, [^2] See Bender and Koller, (2020), Piantadosi and Hill, (2022), and many others. and we do not hope to settle it here. Instead we will grant that the internal states of the models are representational, and then focus on a different problem tied to models having beliefs about the world. This problem is not as widely discussed and arises even if we assume, as we now do, that models have representational states about the world. The worry is that none of these representational states are belief states since they do not meet the minimal requirements for a state being a belief state. This problem is the focus of this section.
@@
Not all representational states have this standard of correctness; belief is special in this regard. To illustrate this point, consider the act of imagining. Imagining is a representational state just like belief, but it does not have truth as its standard of correctness. If I imagine that I am the head chef at a French restaurant, then simply because this is false, there is nothing wrong with this imagining. But if I believe that I am the head chef at a French restaurant, then this belief is defective simply because it is false. There might be all kinds of benefits that come with having this false belief, like increased confidence in the kitchen, but as a state of belief it is defective simply because it is false. This issue is often summarized in the slogan that truth is the norm of belief or that belief aims at the truth: it is in the nature of this internal state that truth is part of its standard of correctness. [^3] The connection of belief to truth is widely discussed in the philosophical literature. See, for example, Williams, (1973), Railton, (1994),Velleman, (2000).
@@ 3 Coherence norms for language models
Belief in humans is subject to the norms of rationality. We are not always rational, but we ought to be, which is to say: we are under the norms of rationality, even if we don’t always live up to those requirements. Rationality itself is complex, and the requirement to be rational can be analyzed as a series of distinct norms. One core part of these norms are coherence norms, which concern how the internal states of a system ought to relate to each other. Other norms of rationality concern the proper response to reasons or to evidence. How the different parts of rationality relate to each other is itself controversial, but we will simply focus on coherence norms in the following. [^4] See Kolodny, (2005), Kiesewetter, (2017), Worsnip, (2021) and many others for the debate about whether coherence norms are derivative on norms tied to responding to reasons or whether they are an independent group of norms of rationality.
@@ 3.2 Defining credence for language models
On the one hand, it is clear that ‘I am sure that $p$’ is a form of assenting to $p$, and that prima facie speaks in favor of including it as an assent sequence. But doing so would have a negative impact in a related area. When a model epistemically marks its response, then this marking should correspond to its confidence in the answer, and thus it should correspond to the credence it has in this being the right answer. So, when a model responds ‘I am certain that $p$’ then this is the proper response only if the credence in $p$ itself is very high. One should only say that one is certain that $p$ if one has high credence in $p$. In fact, this connection can be seen as a further coherence norm for rational thinkers. If the strength of my epistemic marking diverges from my credence, then I am arguably violating a coherence norms. If, however, we include epistemic markers in the list of assent and dissent sequences themselves, then the credence of a model in $p$ will in part be determined by its epistemic markings of $p$. But these issues should at first be kept separate. It is one thing how strongly the model believes that $p$, and another whether a particular epistemic marking is appropriate for that model. It is thus overall best to exclude such epistemically marked sentence from the list of assent and dissent sequences. Maybe it is best to think of ‘I am quite sure that $p$’ as containing two parts, made more explicit by its reformation as ‘$p$ and I am quite sure of it’. The first conjunct is a proper assent to $p$, while the second one is not, and only a comment on the model’s own epistemic state. We should only count the first conjunct as an assent sequence, but not the combination of both as a further one.
Whether or not present day models’ credences match their epistemic markings is an empirical questions. If they do they are epistemically better functioning than if they don’t. To determine whether there indeed is this correlation is subject to ongoing work. See also Zhou et al., (2024) for more on models reporting on their own epistemic state.
@@ 3.3 The argument for probabilism
When focusing on synchronic coherence norms, this question quickly turns into the question whether the credence function should satisfying the axioms of probability theory and thus be a probability function. In the philosophical literature the view that a credence function should be a probability function is called probabilism. [^5] See Joyce, (1998), Teitelbaum, (2023). Assuming credences for language models as spelled out above, does probabilism apply to them as well? Are language models under the coherence norm that their credence function satisfies the axioms of probability theory? For example, are they under the norm that $cr(p)=1-cr(\neg p)$?
@@
However, things are again different for the model that is fine-tuned on truthfulness. A model that aims at true beliefs is subject to the well-known accuracy arguments for probabilism. See Joyce, (1998), Teitelbaum, (2023). The main idea of these arguments is as follow: if the goal is truth, then credences should be as close to the truth as possible: one’s credence in a true proposition should be close to 1, and one’s credence in a false proposition should be close to 0. However close one comes to this ideal can be seen as the accuracy of one’s credences. A standard measure of accuracy in this sense is the Brier score, basically the Euclidean distance between one’s credences in some propositions and their actual truth values (taken as 1 for true, and 0 for false) seen as points in a Euclidean space of the dimension that corresponds to the number of propositions, but without taking a square root. So, the Brier score for a credence function $cr$ from a finite set of propositions $p_{1},...p_{n}$ to [0,1] is defined as
@@
whereby $tv(p)=1$ if p is true and 0 if it is false. [^6] The Brier score is sometimes also given as the mean squared error, i.e. dividing the above by the dimension $n$. This difference does not matter for us here. The thought is that a credence function is better if its Brier score is lower, since one aims at the truth, and a lower Brier score of one’s credence function means that one’s credences are closer to how they should be. In Joyce, (1998), James Joyce showed that if one’s credence function does not meet the axioms of probability theory, then there is a different credence function that has a lower Brier score no matter what the world is like and which meets the axioms of probability theory. However, if one’s credence function does meet the axioms of probability theory, then this is not always the case. This argument is taken to support probabilism: A credence function that isn’t a probability function can always be dominated by one that is. Thus such a credence function isn’t what it should be, given that one aims at the truth. And thus probabilism is a rational norm that applies to ones credences: one’s credence function ought to be a probability function. [^7] For further discussion of Joyce’s argument, see Teitelbaum, (2023) volume 2, pages 338-71. This argument applies to the credence functions of language models just as it applies to the credence functions of human beings. As long as we aim at the truth and at accuracy, i.e. lower Brier scores, our credence functions ought to be probability functions, or so the norms of rationality demand.
@@ 4 Belief revision norms for language models
Language models nonetheless can change their beliefs: they can be fine-tuned to adopt a particular belief or whole set of beliefs, and they can be directly edited to change a particular belief. [^8] Zhu et al., (2020); De Cao et al., (2021); Dai et al., (2021); Hase et al., (2021) But such change of belief of the model is not a rational process for the model itself. It is simply an external modification of the model, and the question of rationality does not directly apply to it from the point of view of the model. It is a change that happens to it from an external source, not a direct response to rational forces by the model itself.
@@ 5 Living up to the normative requirements
The question what norms apply to language models is one that can be discussed rather generally. One can argue, as we did above, that particular classes of models are under particular norms of rationality, while others are not. But the question whether a particular model lives up to the norms that apply to it is one that is specific to each model. It depends on how well the model is doing what it is supposed to do, and that depends on the details of the model, and can’t be determined in general. Nonetheless, there is a general question in the neighborhood, which is one about the effectiveness of rational requirements on models in general. Our human minds are under rational norms and somehow these norms are generally effective. Although we are often irrational, it is not the case that rational norms have no effect on us at all. They often push us in the right direction. How rational and irrational human beings are is a widely studied question. [^9] Tversky and Kahneman, (1974) Our concern here is not human beings, but language models. And to determine how rational they are faces a special additional difficulty: Language models face a conflict of norms that does not seem to carry over to human rationality.
@@
To solve this problem, and to properly test how effective rational norms are in language models, we need to control one variable to test for the other. In particular, we need to control the incoherence of the training data in pre-training to test for rationality in the fine-tuned models. If we knew, say, that the model was only pre-trained on logically consistent training data, then any inconsistency in its beliefs is a result of a rational failure. And if we knew that if the training data was only slightly incoherent, then the degree of incoherence might correspond to a degree of irrationality. But, of course, such experiments are not possible with present language models, since they are trained on a large corpus of text that is not controlled for coherence, and in fact is known to be vastly incoherent. And what is worse, it is known to be hard to test any large body of natural language text for logical coherence. Just to test for Boolean satisfiability is an NP-complete problem (Cook,, 1971). It is likely that pre-training work involving synthetic corpora will be required to determine whether language models produce inconsistencies as a result of adherence to their training data or a failure of rationality.

### v2

@@ 1 Introduction
To get clearer on our main topic, it is important to first note that the term ‘rational’ is used in different ways, and depending on how one understands it, our main question either becomes a trivial one or a more substantial one. On the one hand, ‘rational’ is often used simply as a term of approval, where an AI is rational just in case it does what it is supposed to do. See Russell (1997). On such an understanding, intelligent systems are rational, as long as they work properly. But this conception of being rational does not consider how the system does what it is supposed to do, and what internal states of the system contribute to its success. A different sense of ‘rational’ focuses on these internal states instead, not just on the overall behavior of the system. This is how we often evaluate human beings as rational. Here, we do not simply evaluate their behavior, but we focus on their internal states and evaluate them. We consider their beliefs or reasoning or actions as rational, and we thereby evaluate them as having arisen in the proper way from the proper internal, representational states. We can call the former, thinner notion of being rational the deflationary notion, and the latter the representational notion, since it concerns internal representational states.
@@
In the literature on language models, there is much discussion about whether language models have the ability to reason, which is a cross-temporal rational ability. See Huang and Chang (2023). Reasoning involves a sequence of steps that are properly rationally supported by earlier stages in the process. But the question whether language models engage in reasoning is hard to properly address without first settling whether language models are rational at all and whether norms of rationality even apply to them. The simplest case for making progress on that basic question is to focus first on synchronic coherence norms and on belief states governed by such norms.
Thus to ask whether language models are subject to norms of (representational) rationality like coherence norms involves asking firstly whether they have the proper internal representational states like belief states that are in principle subject to such norms, and secondly whether the models with those states indeed are subject to such norms. To get clear on this issue it is important to distinguish two ways in which a language model can fail to be rational. The first concerns whether a language model is under the norms of rationality in the first place. Here ‘rational’ contrasts with being arational: not being subject to the norms of rationality. The second concerns whether the model lives up to the norms of rationality that apply to it. Here ‘rational’ contrasts with being irrational: not living up the norms of rationality that apply. A toaster is not rational in the first sense, while a person with contradictory beliefs is not rational in the second sense. An arational toaster does not fail to live up the standards of rationality, since those standards do not apply to it at all, while an irrational person does fail this way. We need to ask both of these questions of language models: do the norms of rationality apply to them at all, and if so, do the models live up to those norms. In this paper we will focus mostly on the first, more normative or philosophical question. In a different paper, Hase et al. (2024), we consider the empirical question, the results of which will also be discussed below.
@@ 2 Belief in language models
The easiest way to argue that language models have beliefs is via the connection of belief to knowledge. Language models seem to know a lot about the world. Although they might make things up and get things wrong, they nonetheless often reliably produce correct answers to many questions about the world. For example, many large language models can correctly answer the question of what the capital of France is and thus are naturally taken to know that the capital of France is Paris. But knowledge is just a special kind of belief: a belief that is true and especially well supported. The details of the connection between knowledge and belief have long been controversial in the philosophical literature. In particular, it is controversial what else is required besides a state being a belief state that has a true content for this state to be knowledge. Nonetheless, the general connection between knowledge and belief relied upon here is almost universally accepted: to know is to believe. Thus since language models know about the world, they have beliefs about the world. [^1] For a discussion of the relationship between belief and knowledge, see Ichikawa and Steup (2018).
@@
To have beliefs about the world requires an internal state that is a state of believing, and simply carrying information alone is not enough for this. The question is whether language models have internal states which are states of belief. There are two main worries about this, one is very familiar and one less so. The first concerns whether the internal states of language models are representational states at all. Do any internal states of the model represent the world which created the text on which they are trained? Or does the model merely reliably predict text without any regard to the semantic properties of the text they produce? This issue is widely discussed, [^2] See Bender and Koller (2020), Piantadosi and Hill (2022), and many others. and we do not hope to settle it here. Instead we will grant that the internal states of the models are representational, and then focus on a different problem tied to models having beliefs about the world. This problem is not as widely discussed and arises even if we assume, as we now do, that models have representational states about the world. The worry is that none of these representational states are belief states since they do not meet the minimal requirements for a state being a belief state. This problem is the focus of this section.
@@
Not all representational states have this standard of correctness; belief is special in this regard. To illustrate this point, consider the act of imagining. Imagining is a representational state just like belief, but it does not have truth as its standard of correctness. If I imagine that I am the head chef at a French restaurant, then simply because this is false, there is nothing wrong with this imagining. But if I believe that I am the head chef at a French restaurant, then this belief is defective simply because it is false. There might be all kinds of benefits that come with having this false belief, like increased confidence in the kitchen, but as a state of belief it is defective simply because it is false. This issue is often summarized in the slogan that truth is the norm of belief or that belief aims at the truth: it is in the nature of this internal state that truth is part of its standard of correctness. [^3] The connection of belief to truth is widely discussed in the philosophical literature. See, for example, Williams (1973), Railton (1994),Velleman (2000).
@@ 3 Coherence norms for language models
Belief in humans is subject to the norms of rationality. We are not always rational, but we ought to be, which is to say: we are under the norms of rationality, even if we don’t always live up to those requirements. Rationality itself is complex, and the requirement to be rational can be analyzed as a series of distinct norms. One core part of these norms are coherence norms, which concern how the internal states of a system ought to relate to each other. Other norms of rationality concern the proper response to reasons or to evidence. How the different parts of rationality relate to each other is itself controversial, but we will simply focus on coherence norms in the following. [^4] See Kolodny (2005), Kiesewetter (2017), Worsnip (2021) and many others for the debate about whether coherence norms are derivative on norms tied to responding to reasons or whether they are an independent group of norms of rationality.
@@ 3.2 Defining credence for language models
On the one hand, it is clear that ‘I am sure that $p$’ is a form of assenting to $p$, and that prima facie speaks in favor of including it as an assent sequence. But doing so would have a negative impact in a related area. When a model epistemically marks its response, then this marking should correspond to its confidence in the answer, and thus it should correspond to the credence it has in this being the right answer. So, when a model responds ‘I am certain that $p$’ then this is the proper response only if the credence in $p$ itself is very high. One should only say that one is certain that $p$ if one has high credence in $p$. In fact, this connection can be seen as a further coherence norm for rational thinkers. If the strength of my epistemic marking diverges from my credence, then I am arguably violating a coherence norm. If, however, we include epistemic markers in the list of assent and dissent sequences themselves, then the credence of a model in $p$ will in part be determined by its epistemic markings of $p$. But these issues should at first be kept separate. It is one thing how strongly the model believes that $p$, and another whether a particular epistemic marking is appropriate for that model. It is thus overall best to exclude such epistemically marked sentence from the list of assent and dissent sequences. Maybe it is best to think of ‘I am quite sure that $p$’ as containing two parts, made more explicit by its reformation as ‘$p$ and I am quite sure of it’. The first conjunct is a proper assent to $p$, while the second one is not, and only a comment on the model’s own epistemic state. We should only count the first conjunct as an assent sequence, but not the combination of both as a further one.
Whether or not present day models’ credences match their epistemic markings is an empirical question. If they do they are epistemically better functioning than if they don’t. To determine whether there indeed is this correlation is subject to ongoing work. See also Zhou et al. (2024) for more on models reporting on their own epistemic state.
@@ 3.3 The argument for probabilism
When focusing on synchronic coherence norms, this question quickly turns into the question whether the credence function should satisfying the axioms of probability theory and thus be a probability function. In the philosophical literature the view that a credence function should be a probability function is called probabilism. [^5] See Joyce (1998), Teitelbaum (2023). Assuming credences for language models as spelled out above, does probabilism apply to them as well? Are language models under the coherence norm that their credence function satisfies the axioms of probability theory? For example, are they under the norm that $cr(p)=1-cr(\neg p)$?
@@
However, things are again different for the model that is fine-tuned on truthfulness. A model that aims at true beliefs is subject to the well-known accuracy arguments for probabilism. See Joyce (1998), Teitelbaum (2023). The main idea of these arguments is as follow: if the goal is truth, then credences should be as close to the truth as possible: one’s credence in a true proposition should be close to 1, and one’s credence in a false proposition should be close to 0. However close one comes to this ideal can be seen as the accuracy of one’s credences. A standard measure of accuracy in this sense is the Brier score, basically the Euclidean distance between one’s credences in some propositions and their actual truth values (taken as 1 for true, and 0 for false) seen as points in a Euclidean space of the dimension that corresponds to the number of propositions, but without taking a square root. So, the Brier score for a credence function $cr$ from a finite set of propositions $p_{1},...p_{n}$ to [0,1] is defined as
@@
whereby $tv(p)=1$ if p is true and 0 if it is false. [^6] The Brier score is sometimes also given as the mean squared error, i.e. dividing the above by the dimension $n$. This difference does not matter for us here. The thought is that a credence function is better if its Brier score is lower, since one aims at the truth, and a lower Brier score of one’s credence function means that one’s credences are closer to how they should be. In Joyce (1998), James Joyce showed that if one’s credence function does not meet the axioms of probability theory, then there is a different credence function that has a lower Brier score no matter what the world is like and which meets the axioms of probability theory. However, if one’s credence function does meet the axioms of probability theory, then this is not always the case. This argument is taken to support probabilism: A credence function that isn’t a probability function can always be dominated by one that is. Thus such a credence function isn’t what it should be, given that one aims at the truth. And thus probabilism is a rational norm that applies to ones credences: one’s credence function ought to be a probability function. [^7] For further discussion of Joyce’s argument, see Teitelbaum (2023) volume 2, pages 338-71. This argument applies to the credence functions of language models just as it applies to the credence functions of human beings. As long as we aim at the truth and at accuracy, i.e. lower Brier scores, our credence functions ought to be probability functions, or so the norms of rationality demand.
@@ 4 Belief revision norms for language models
Language models nonetheless can change their beliefs: they can be fine-tuned to adopt a particular belief or whole set of beliefs, and they can be directly edited to change a particular belief. [^8] Zhu et al. (2020); De Cao et al. (2021); Dai et al. (2021); Hase et al. (2021) But such change of belief of the model is not a rational process for the model itself. It is simply an external modification of the model, and the question of rationality does not directly apply to it from the point of view of the model. It is a change that happens to it from an external source, not a direct response to rational forces by the model itself.
@@ 5 Living up to the normative requirements
The question what norms apply to language models is one that can be discussed rather generally. One can argue, as we did above, that particular classes of models are under particular norms of rationality, while others are not. But the question whether a particular model lives up to the norms that apply to it is one that is specific to each model. It depends on how well the model is doing what it is supposed to do, and that depends on the details of the model, and can’t be determined in general. Nonetheless, there is a general question in the neighborhood, which is one about the effectiveness of rational requirements on models in general. Our human minds are under rational norms and somehow these norms are generally effective. Although we are often irrational, it is not the case that rational norms have no effect on us at all. They often push us in the right direction. How rational and irrational human beings are is a widely studied question. [^9] Tversky and Kahneman (1974) Our concern here is not human beings, but language models. And to determine how rational they are faces a special additional difficulty: Language models face a conflict of norms that does not seem to carry over to human rationality.
@@
To solve this problem, and to properly test how effective rational norms are in language models, we need to control one variable to test for the other. In particular, we need to control the incoherence of the training data in pre-training to test for rationality in the fine-tuned models. If we knew, say, that the model was only pre-trained on logically consistent training data, then any inconsistency in its beliefs is a result of a rational failure. And if we knew that if the training data was only slightly incoherent, then the degree of incoherence might correspond to a degree of irrationality. But, of course, such experiments are not possible with present language models, since they are trained on a large corpus of text that is not controlled for coherence, and in fact is known to be vastly incoherent. And what is worse, it is known to be hard to test any large body of natural language text for logical coherence. Just to test for Boolean satisfiability is an NP-complete problem (Cook, 1971). It is likely that pre-training work involving synthetic corpora will be required to determine whether language models produce inconsistencies as a result of adherence to their training data or a failure of rationality.
In related empirical work, Hase et al. (2024), we hope to make progress on these questions in just this way: by testing rationality in a new language model that is pre-trained from scratch on a novel semi-synthetic dataset controlled for coherence. In particular, we hope to assess probabilistic coherence in models pre-trained on datasets that are incoherent in a measurable way. To do this properly, we need to work with semi-synthetic data whose coherence is controllable in a precise way. In a nutshell, it appears that the model’s internal probabilities track the degree of incoherence in the training data, which largely speaks in favor of rational norms being effective. However, when it comes to model updating and belief revision, present methods fall far short of an ideal Bayesian agent. Again, a full discussion of this empirical work is in Hase et al. (2024).